Let P (x): x2/2 =x


find the following and identify their truth values

1. (P) 1

2. (P) 2

3.∀n, P(n)

4.∃n, P(n)


1
Expert's answer
2021-03-31T17:04:12-0400
  1. "\displaystyle P(1): \frac{1^2}{2}= \frac{1}{2} \neq 1 \Rightarrow \text{False}"
  2. "\displaystyle P(2): \frac{2^2}{2}= 2 =2 \Rightarrow \text{True}"
  3. in verbs: for all n P(n). This is false, because we can give a contrexample (like P(1)).
  4. in verbs: exists such a n that P(n). This is true, because as we saw for P(2) such n=2 indeed exists.

Answer: 1.F, 2. T, 3. F, 4. T.